OpenStudy (anonymous):

solve the equation:1.045^t=2

OpenStudy (dumbcow):

take log of both sides use following log property: log(x^n) = nlog(x)

OpenStudy (anonymous):

y (1.045)^t = 2 log(1.045)2 = t t

OpenStudy (anonymous):

take log of both sidesuse following log property: log(x^n) = nlog(x)